OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1.

jimthompson5910 (jim_thompson5910):

5(2x+1)=12x-1 10x+5=12x-1 10x=12x-1-5 10x-12x=-1-5 -2x=-1-5 -2x=-6 x=-6/(-2) x=3